OpenStudy (anonymous):

What is the equation of the line expressed in slope-intercept form? A. y = –3x – 1 B. y = 3x + 2 C. y = 2x + 3 D. y = –2x + 3

OpenStudy (anonymous):

@pooja195 Can you help?

pooja195 (pooja195):

y=mx+b <--slope intercept form

pooja195 (pooja195):

slope=rise/run any ideas on what the slope if?

OpenStudy (anonymous):

nope

pooja195 (pooja195):

Ok, Look here http://prntscr.com/9pel2d how much are we rising?

OpenStudy (anonymous):

4

pooja195 (pooja195):

Count the units that the pink line is going up

OpenStudy (anonymous):

3

OpenStudy (anonymous):

it goes up 2 units

pooja195 (pooja195):

good 2 is right how much are we going over?

OpenStudy (anonymous):

a third of the way?

pooja195 (pooja195):

No1 unit to the right so 2/1=?

OpenStudy (anonymous):

2

pooja195 (pooja195):

good so slope =2 now to find the y intercept.... Where is the line hitting on the y axis?

OpenStudy (anonymous):

2?

pooja195 (pooja195):

Good now put it into an equation... y=mx+b m=slope b=y intercept

OpenStudy (cheyennecoffelt):

so slope and y-intercept are the same? y=2x+2?

pooja195 (pooja195):

And you just gave the answer ,-,

OpenStudy (cheyennecoffelt):

but it isn't up there

OpenStudy (cheyennecoffelt):

It isn't A, B, C, or D

pooja195 (pooja195):

My bad :P If y intercept is 2 there is only one answer that has +2 as the y intercept (i had the slope wrong)

OpenStudy (anonymous):

I'm back sorry I was in the bathroom